You wish to install a solar panel that will run at least five 15-W LED lightbulbs, an LCD TV (300 W), and a microwave (1500 W). What is the area the panel should have if the average sunlight incident on a photoelectric solar collector on a roof for the 8-h time interval is 800 W/m2 and the radiant energy is converted to electricity with an efficiency of 23 %?
Answer #1

Total power to be delivered to run the electrical appliances is given by, P = (5*15+300+1500) = 1875 W

Input energy available per unit area = (800*8*3600) J/m2

= 23040000 J/m2

Thus, output enrgy = efficiency*input energy. = (23/100 * 23040000) = 5299200 J/m2

So, 1875*8*3600 = 5299200*A

Area required, A = (1875*8*3600/5299200) = 10.19 m2